Solve the equation and check.
A. -3(x-3)=2x

Answer by Maths68(1474)

-3(x-3)=2x
-3x+9=2x
-3x-2x=-9
-5x=-9
-5x/-5=-9/-5
x=9/5


Check
=====
-3(x-3)=2x
Put x=9/5
-3(9/5-3)=2(9/5)
-27/5+9=18/5
(-27+45)/5=18/5
18/5=18/5